# A discussion about violin reduction: geometric analysis of contour lines and channel of minima

### v1

@@ Appendix 0.A List of instruments in the corpus
This table lists the instruments studied in this research, classified by their inventory number at the Musical Instruments Museum (MIM), Brussels. The hyperlink on each number refers to the official MIM documentation and provides additional information. The question marks in the table indicate missing data or uncertainties. Finally, four instruments stand out slightly from the rest of the corpus: two tenor violins and two bass violins. These are names and sizes used before the standardisation of instruments, around 1750. The tenor violins are slightly larger than violas, and the bass violins are slightly larger than cellos. For ease of reference, the 2836 has been compared with the same scale as violins and violas in Sections 4 and 5, while the 2853 and the two bass violins have been associated with cellos. It should be noted that although the two tenor violins share the same terminology, the two instruments are quite different. The 2853 has much larger ribs, and is thought to have been used as a child’s cello, which explains its comparison with cellos.
@@ 0.A.0.1 Acknowledgements
We thank Pr. Sandra Soares Frazao and her research team for sharing their best practices in photogrammetry. We would also like to thank Joris De Valck of the MIM (Brussels), Mirte Maes and Karel Moens of the Museum Vleeshuis (Antwerp), and Jan Strick of the Maison Bernard (Brussels). Their expertise on early instruments strengthened the analysis and discussion of our results.

### v2

@@ Appendix 0.A List of Instruments in the Corpus
This table lists the instruments studied in this research, classified by their inventory number at the Musical Instruments Museum (MIM), Brussels. The hyperlink on each number refers to the official MIM documentation and provides additional information. The question marks in the table indicate missing data or uncertainties. Finally, four instruments stand out slightly from the rest of the corpus: two tenor violins and two bass violins. These are names and sizes used before the standardisation of instruments, around 1750. The tenor violins are slightly larger than violas, and the bass violins are slightly larger than cellos. For ease of reference, the 2836 has been compared with the same scale as violins and violas in Sections 4 and 5, while the 2853 and the two bass violins have been associated with cellos. It should be noted that although the two tenor violins share the same terminology, the two instruments are quite different. The 2853 has much larger ribs, and is thought to have been used as a child’s cello, which explains its comparison with cellos (Table 1).
@@ 0.A.0.1 Acknowledgements
We thank Pr. Sandra Soares Frazao and her research team for sharing their best practices in photogrammetry. We would also like to thank Joris De Valck of the MIM (Brussels), Mirte Maes and Karel Moens of the Museum Vleeshuis (Antwerp) and Jan Strick of the Maison Bernard (Brussels). Their expertise on early instruments strengthened the analysis and discussion of our results. We thank Jean-Philippe Echard of the Musée de la Musique - Philharmonie de Paris for his careful reading and insightful comments on a previous version of this article. Finally, we would also like to thank Iona Thys for the photographic acquisition of the corpus of instruments.